# Estimating energy levels from lattice QCD correlation functions using a transfer matrix formalism

### v1

@@ Appendix B Numerical comparison between Lanczos and TGEVP methods
Below we compare the numerical results obtained with the Lanczos method, as described in [1], and the TGEVP method, as proposed here. We follow a similar implementation of Oblique Lanczos method (Lanczos A) as in Ref. [1]. We also use another method (Lanczos B) which is similar to Ref. [1], except differences in the procedure of filtering and sorting the eigenvalues. We outline the implementation below.
@@
    We first resample our data to obtain $N_{b}$ bootstrap samples and perform $M$ iterations of Oblique Lanczos for each, obtaining $\lambda^{m,b}_{n}$. The index $1\leq M$ and $1\leq b\leq N_{b}$ correspond to Lanczos iterations and the bootstrap sample being used respectively, while $n$ enumerates the computed eigenvalues.
@@
In Fig. 6, we compare the eigenvalues obtained using the Lanczos and TGEVP methods for the $\eta_{c}$ and $\Xi_{cc}$, corresponding to the dateset shown in the respective top panes of Figs. 1 and 2. This comparison suggests the numerical equivalence of the two methods.
@@
Appendix C Additional results
In addition to the results presented in the main section, here we present more such results for various hadrons including their excited states. In Fig. 7, we show the first two eigenvalues as determined employing TGEVP method. These dataset corresponds to 1S-charmonia states, where the left panel is for $0^{-}$ and the right one corresponds to $1^{-}$ states. The regular effective mass, $m_{eff}(t)$ is also shown (blue circles) to compare that with the lowest eigenvalue (red circles). In addition to the lowest eigenvalue we find the second eigenvalue (green circles) is also stable across iterations, and one can get a good estimation of the first excited state. These results are obtained from the two point correlation functions computed with overlap valence quarks on the $N_{f}=2+1+1$ HISQ lattice ensembles with the specifications, $L^{3}\times T$ (lattice spacing $a$ in fm) $\equiv$ $32^{3}\times 96~{}(0.0888)$ (top), $48^{3}\times 144~{}(0.0582)$ (middle), and $64^{3}\times 192~{}(0.0441)$ (bottom) [27, 28], and was discussed in Refs. [8, 29, 30] . In Fig. 8, we present similar results using the HISQ action for valence quarks on the same set of lattices. We show the effective mass and the first two/three TGEVP-eigenvalues of charmonia states. Left panel represents the $J^{P}~{}\equiv~{}0^{-}$ state while the right one corresponds to $1^{-}$ and $1^{+}$ states. In this case $1^{-}$ requires special attention due to oscillating contributions in the temporal direction. To address that, we employ the smoothed effective mass technique for the ground state of $J/\psi$, as described in [31]. From the non-oscillating part of the correletors we extract the lowest two eigenvalues which correspond to the masses of $J/\psi(1S)$ and $J/\psi(2S)$ states. The lowest eigenvalue corresponding to the oscillating part of the correlator is also extracted and identified that with the mass of $\chi_{c1}(1P)$ state. For the eigenvalue of the oscillating state, only the negative eigenvalues, $\lambda_{n}^{m,b}$, of Eq. (9) are retained and then we follow the same analysis procedure as in Sec. (II.1).
Figure 7: These plots represent the regular effective mass and the first two TGEVP-eigenvalues corresponding to 1S-charmona states. Left panel is for $0^{-}$ and the right one corresponds to $1^{-}$. These results are obtained with overlap valence quarks on the $N_{f}=2+1+1$ HISQ lattice ensembles with the specifications, $L^{3}\times T$ (lattice spacing $a$ in fm), $n_{meas}$equiv Top: $32^{3}\times 96~{}(0.0888)$, 169, Middle: $48^{3}\times 144~{}(0.0582)$, 180, and Bottom: $64^{3}\times 192~{}(0.0441)$, 142.
@@ Appendix C Additional results
Based on these results on the lowest two/three eigenstates we then calculate various energy splittings between them, for example, hyperfine splitting in 1S, along with 2S--1S and 1P--1S splittings in charmonia. Since we could extract these values at multiple lattice spacings we then extrapolate them to the continuum limit. Since both the overlap and HISQ actions do not have $\mathcal{O}(n\times ma)$ errors ($n$ odd integers, $ma$ being quark mass in lattice unit), we extrapolate these energy splittings with a fit ansatz, $f(a)=c_{1}+c_{2}a^{2}+c_{3}a^{4}$. We show the continuum extrapolated results in Figs. [9,10, 11]. Fig. 9 represents the hyperfine splitting for 1S-charmonia obtained with both overlap (red square) and HISQ (blue square) actions. The continuum extrapolated results are shown by the red and green star-symbols respectively, while the experimental value is shown by the magenta square. In Fig. 10, we present the 2S --1S splitting with overlap results at the top and the HISQ results at the bottom pane. Here again, we use the same fit form and extrapolations are shown with the same symbols as above. In Fig. 11, we show the 1P--1S splitting.
@@
Figure 10: Continuum extrapolation of the charmonia energy splittings, using the lowest two eigenvalues, obtained by TGEVP method as shown in 7. Top: results with overlap valence quarks, Bottom: same with HISQ valence quarks. The star symbols represent the continuum extrapolated values while solid squares are their experimental values.
@@
In Fig. 12, we present the second eigenvalue corresponding to the nucleon correlator that was shown in Fig. 3. Here we plot them together. Note that this second energy level may correspond to the Roper resonance. It needs to be seen if this second energy level can be extracted robustly at the lighter pion mass, in the presence of $N\pi$ $N\pi\pi$ states, leading to the physical Roper resonance.
@@
Figure 12: Second energy level together with the lowest level corresponding to the nucleon correlators shown earlier in Fig. 3. Note the second level may correspond to the Roper resonance.
In Fig. 13, we present results for nuclei (similar to Fig. 4) but for the case with $m_{u}=m_{d}=m_{c}$. Here we show results for 7Li also. Again for each case we find the TGEVP eigenvalues are much more reliable compared to the fluctuating effective mass. For the cases of 4He and 7Li, we observe that the lowest GEVP eigenvalue plateau is somewhat lower than the probable effective mass plateau. That leads to the following question: Whether this is because of the presence of a large number of close-by states, which becomes denser at the higher quark masses, TGEVP fails or the effective mass in this asymmetric source-sink setup becomes quite unreliable? This is an interesting observation and needs to be investigated when more statistics become available.
Figure 13: Lowest energy eigenvalues of a few light nuclei, up to 7Li, computed from their two-point functions with the quark masses $m_{u}=m_{d}=m_{c}$. Respective effective masses are also shown for comparison.
Appendix D Detecting peaks from the histograms of eigenvalues
Due to presence of statistical noises spurious eigenvalues appear and that skew their distribution. For a physical state the eigenvalue distribution should have a peak with a width. Hence, the primary objective in detecting peaks from the histograms of eigenvalues is to identify the centers and widths of all peaks present in the histogram. To mitigate the influence of false peaks caused by statistical fluctuations in the data, we begin by smoothing the histogram counts using a Gaussian filter. In a smoothed histogram it becomes easier to detect all peaks that exceed a predefined threshold count. In Fig. 14 we show two such peaks after Gaussian filtering of the original data. This data corresponds to the two-point correlation function of the $1^{-}$ states in charmonia and represents the two states showed in the bottom right pane of Fig. 7, $J/\psi(1S)$ and $J/\psi(2S)$.
Figure 14: Histogram plots showing two peaks as determined after Gaussian filtering as described in the text. This data corresponds to the two-point correlation function of the $1^{-}$ states in charmonia and represents the two states showed in the bottom right pane of Fig. 7, $J/\psi(1S)$ and $J/\psi(2S)$.
Appendix E Correlation in TGEVP eigenvalues
@@
The eigenvalues obtained in different iterations of the TGEVP method are inherently correlated, as each iteration builds upon the results of the previous one through random sampling. To assess the extent of this correlation, we compute the associcated correlation matrix using the bootstrap method from the eigenvalues that fall within the 68th percentile interval. The first construct the covariance matrix as,
@@
We present here this correlation matrix in a color-coded format for a few eigenvalues presented in the main text. In Fig. 15, we show a plot showing correlation matrix corresponding to the excited state of $1^{-}$ state in charmonia, This corresponds to the second eigenvalue of the top right plot of Fig. 7. As can be seen that correlation is somewhat larger compared to that of the lowest eigenvalues shown in Fig. 5. The higher eigenvalues corresponding to other correlation functions also show the same pattern of higher correlation than that present in the lowest eigenvalue. In Fig. 16 we depict the correlation matrix plot of the first two eigenvalues of the nucleon two-point functions corresponding to the two states shown in Fig. 12. Here again, the second state, which may correspond to the Roper resonance, shows more correlation than the ground state.
@@
In Fig. 17 we show correlation matrix for the light nuclei. The top pane corresponds to deuteron correlators at $m_{u}=m_{d}=m_{s}$ (left) and $m_{u}=m_{d}=m_{c}$ (right), respectively. The bottom pane represents the same for 4He. It clearly shows that correlation increases at the lighter quark mass and also increases with the atomic number. The eigenvalues corresponding to 4He correlators at the strange quark mass are correlated to a large extent, and increases further as the quark mass become lighter.
@@
This analysis leads us to the following observations: 1. If the statistics is high and the effective mass is reasonably good we do not find much correlation of eigenvalues across iterations. In that case fitting these eigenvalues with a constant term will be reliable and provide better estimate compared to exponential fitting. This could be a good choice in multi-operator GEVP setup, that is merging of GEVP and TGEVP together. We will explore that in future 2. For the higher eigenvalues, correlations are larger than that for the case of lowest eigenvalue. With higher statistics we find the correlation in the higher eigenvalues also decreases. 3. For multi-nucleons we find correlations to be higher than the single nucleon. This is a concern and needs to be addressed properly in a multinucleon calculation, particularly considering the presence of dense spectrum. It remains to be seen how this correlation evolves with higher statistics in these systems.
@@
Figure 15: Correlation matrix plot corresponding to the second eigenvalue of the two-point correlation function of $1^{-}$ state in charmonis ($J/\psi(2S)$ state.)

### v2

@@ V.2 Numerical comparison between Lanczos and TGEVP methods
The analytical equivalence between the Lanczos and TGEVP methods are shown in the Appendix A. Here we compare the numerical results obtained with the Lanczos method, as described in Ref. [1], and the TGEVP method, as proposed here. We follow a similar implementation of Oblique Lanczos method (Lanczos A) as in Ref. [1]. We also use another method (Lanczos B) which is similar to Ref. [1], except differences in the procedure of filtering and sorting the eigenvalues. We outline the implementation below.
@@
    We first resample our data to obtain $N_{b}$ bootstrap samples and perform $M$ iterations of Oblique Lanczos for each, yielding $\lambda^{m,b}_{n}$. The index $1\leq M$ and $1\leq b\leq N_{b}$ correspond to Lanczos iterations and the bootstrap sample being used respectively, while $n$ enumerates the computed eigenvalues.
@@
Following the above procedure for Lanczos method, in Fig. 4, we compare the eigenvalues obtained using the Lanczos and TGEVP methods for the same set of correlation functions corresponding to a meson: $\eta_{c}$ (top), and a baryon: $\Xi_{cc}$ (bottom). Results for Lanczos A and B are both shown. One can clearly observe the numerical equivalence of the two methods, though Lanczos A, where our filtration procedure has not been not employed show fluctuations on some iterations for both the cases.
@@
Having compared the TGEVP method with other existing approaches—and demonstrated the robustness and efficiency of our approach in extracting energy levels—we now present additional results obtained using our proposed method. In Figs. (5, 6, 7, 8, 9, 10, 11, 12) we show our results for a variety of hadrons, both for mesons and baryons. In each of these figures we present the extracted lowest energy levels as estimated by the above method, along with the standard effective mass, defined as $m_{eff}(t)={\mathrm{log}}[C(t)/C(t+1)]$. In each case, throughout this article from here, blue circles represent the regular effective mass $m_{eff}(t)$, and the red circles represent the lowest eigenvalues as computed over the iterations. The magenta band represents the correlated fit of the two-point function using a single exponential, illustrating the fit range and the associated $1\sigma$ error. The faded-green band corresponds to the eigenvalues, fitted with a constant term over multiple iterations, also using a correlated fit. This comparison highlights the stability and error behavior of the TGEVP eigenvalues over regular effective mass approach. Below we discussed each of the cases. A large set of other results, including the first excited states and nuclei, are shown in Appendix C.
V.3 Case I: Charmonia
We first present the results for the 1S-charmonia in Fig. 5. The results corresponds to the two-point correlation functions of $0^{-}$ (top) and $1^{-}$ (bottom) states of charmonia, as computed using overlap valence quarks with a HISQ lattice ensemble ($48^{3}\times 144$, $a=0.0582$ fm, $n_{meas}=180$). As can be observed, while the effective mass is fluctuating over time-slices, the lowest eigenvalue estimated over different iterations is very stable. In Fig 20 in Appendix C we show additional similar results for other lattice ensembles along with the first excited states.
Figure 5: Comparison of the results between the standard effective mass of the ground state (blue circles) and the lowest eigenvalue (red circles) corresponding to the two-point correlation functions of the pseudoscalar ($J^{P}\equiv 0^{-}$) (top) and vector ($1^{-}$) (bottom) states of charmonia. Number of measurements ($n_{meas}$) in this case is $180$.
@@ V.3 Case I: Charmonia
Based on these results on the lowest two/three eigenstates we then calculate various energy splittings between them, for example, hyperfine splitting in 1S, along with 2S--1S and 1P--1S splittings in charmonia. Since we could extract these values at multiple lattice spacings we then extrapolate them to the continuum limit. Since both the overlap and HISQ actions do not have $\mathcal{O}(n\times ma)$ errors ($n$ odd integers, $ma$ being quark mass in lattice unit), we extrapolate these energy splittings with a fit ansatz, $f(a)=c_{1}+c_{2}a^{2}+c_{3}a^{4}$. We show the continuum extrapolated results in Figs. [6,7, 8]. Fig. 6 represents the hyperfine splitting for 1S-charmonia obtained with both overlap (red square) and HISQ (blue square) actions. The continuum extrapolated results are shown by the red and green star-symbols respectively, while the experimental value is shown by the magenta square. Note that the HISQ result agrees with that obtained by HPQCD collaboration using the same lattice ensembles [43]. In Fig. 7, we present the 2S --1S splitting with overlap results at the top and the HISQ results at the bottom pane. Here again, we use the same fit form and extrapolations are shown with the same symbols as above. In Fig. 8, we show the 1P--1S splitting.
@@
Figure 7: Continuum extrapolation of the charmonia energy splittings, using the lowest two eigenvalues, obtained by TGEVP method as shown in 20. Top: results with overlap valence quarks, Bottom: same with HISQ valence quarks. The star symbols represent the continuum extrapolated values while solid squares are their experimental values.
@@
V.4 Case II: Heavy Baryons
After discussing mesons, we now move on to application of TGEVP method in extracting energy levels for baryons, which are generally noisier than those of mesons. This makes it particularly important to assess the efficiency of the TGEVP method in such cases. In Figs. 9 and 10, we show lowest energy levels extracted by TGEVP method for the charmed-baryons, $\Xi_{cc}$ (Fig. 9), $\Omega_{c}$ and $\Omega_{cc}$ (Fig. 10), and compare those with their respective effective masses. The results are obtained from their respective two point functions computed on CLS ensembles generated on $N_{f}=2+1$ nonperturbatively improved Wilson fermions [44]. It again very clearly shows the stability of the lowest eigenvalue over the iteration numbers, as opposed to unstable and fluctuating effective mass data. In such cases, the TGEVP results clearly demonstrate superiority over the effective mass method and related single-exponential fits, showing that it can reliably extract the ground state while avoiding the systematic uncertainties associated with the choice of fit ranges in traditional fitting procedures.
Figure 9: Similar to Fig. 5, corresponding to smeared-smeared two-point functions for the doubly charmed baryon $\Xi_{cc}$, computed at two different lattice spacings with number of measurements ($n_{meas}$) $1612$ and $2000$, respectively.
Figure 10: Similar to Fig. 5, corresponding to smeared-smeared two-point functions for the singly and doubly charmed baryon $\Omega_{c}$ (top) and $\Omega_{cc}$ (bottom), with number of measurements ($n_{meas}$) $1009$.
V.5 Case III: Nucleon and Nuclei
In general, the signal-to-noise ratio (SNR) decreases at lighter pion masses, making it especially important to extract energy levels reliably at those quark masses in order to obtain physically meaningful results. In Fig. 11, we show the lowest energy level obtained from a smeared-point two-point function of the nucleon computed at a near physical pion mass. Note that while the results from the two estimations are consistent, the eigenvalues at the larger iterations have smaller errors compared to errors in the effective mass at the larger times.
@@ V.5 Case III: Nucleon and Nuclei
Figure 11: First two energy levels, corresponding to a smeared-point two-point function of nucleon $(n_{meas}=300)$. Note the second level may correspond to the Roper resonance.
In contrast to regular hadrons, the study of nuclei using lattice QCD remains at a relatively early stage of development. One of the main bottlenecks in that is the lack of reliability in extracting the energy levels, even for the ground states of low-lying nuclei [12, 13, 14, 15]. This is primarily due to poor SNR in the two-point correlation functions at lighter quark masses, resulting very noisy and small plateau at a large distance in the effective mass plots. As a result, statistical error together with fit-uncertainties, becomes large making the whole process unreliable. It is thus quite important to check if TGEVP method can be helpful in this. In Fig. 12, we show the effective mass of the ground state and lowest eigenvalue of TGEVP spectrum for nucleon and 2H. The associated two-point functions are obtained with clover valence quarks on the $N_{f}=2+1$ HISQ gauge ensembles ($0.07664^{3}\times 64,~{}a=$ 0.076 fm) at the quark masses, $m_{u}=m_{d}=m_{s}$. Additional results for 2H, 3He, 4He and 7Li are provided in Fig. 22 of Appendix C, where overlap valence quarks on HISQ gauge ensembles are utilized to compute the two point functions ($N_{f}=2+1+1$ HISQ gauge ensembles ($0.058248^{3}\times 144,~{}a=$ 0.0582 fm) at the quark masses, $m_{u}=m_{d}=m_{s}$: left panel, and $m_{u}=m_{d}=m_{c}$: right panel). The method for computing the correlation functions for these nuclei has been provided in Ref. [15]. As can be observed that the effective masses of these nuclei exhibit significant noise at large times, and sometime also from the beginning (Fig. 22), leading to substantial fit-window error in traditional fitting methods. In contrast, the TGEVP eigenvalues remain stable across numerous iterations. With a large number of measurements, it is expected that errors in eigenvalues in each iteration can be reduced. However, it is also necessary to assess the method’s effectiveness, with the availability of higher statistics, for the cases with closely spaced states carrying finite momentum, particularly when dealing with varied source-sink setups in nuclear systems. Another point of concern is correlation over iterations which we will discuss next. Nevertheless, as the results suggest, this method holds excellent promise for extracting reliable energy levels for nuclei in the future.
Figure 12: Comparison of the results between the regular effective mass (green squares) and TGEVP ground state (red circles) corresponding to two-point functions of a Nucleon (top pane) and Deuteron (bottom pane) with the quark masses, $m_{u}=m_{d}=m_{s}$ $(n_{meas}=9600)$.
V.6 Correlation in TGEVP eigenvalues
@@
The eigenvalues obtained at different iterations of the TGEVP method are inherently correlated, as each iteration builds upon the results of the previous one through random sampling. Addressing this correlation mathematically, where data in each step in a matrix are generated randomly from the previous step, is non-trivial. In Refs. [1, 25] a double-bootstraping is done to address that. In this work, we define the bootstrap confidence interval by using the $68$th percentile.
To assess the extent of this correlation, we compute the associated correlation matrix using the bootstrap method from the eigenvalues that fall within the 68th percentile interval. The first construct the covariance matrix as,
@@
We present here this correlation matrix in a color-coded format for some of the above correlators that we analyzed. In Fig. 13, we show such a color-coded correlation-plot for the two-point correlation functions for $\eta_{c}$ that was shown in the top pane of Fig. 5. As observed, the correlation of eigenvalues across iterations is insignificant at the beginning and increases towards the end. Particularly in the fit region (Fig. 5), it remains small enough to justify a correlated fitting with a constant term.
Figure 13: Color-coded correlation matrix showing the correlations between the eigenvalues across iterations. This dataset corresponds to the two-point correlation functions of the lowest $0^{-}$ state ($\eta_{c}$ meson), showed in the top pane of Fig. 5.
@@
In general, extracting higher energy levels is more challenging than extracting the ground state. Therefore, it is crucial to examine the correlations for the excited states to ensure they can be reliably obtained using the TGEVP method. In Fig. 14, we show correlation matrix corresponding to the excited state of $1^{-}$ state in charmonia, This corresponds to the second eigenvalue of the top right plot of Fig. 20. As can be seen that correlation is somewhat larger compared at the fit window to that of the lowest eigenvalues shown in Fig. 5. The higher eigenvalues corresponding to other correlation functions also show the same pattern of higher correlation than that present in the lowest eigenvalue. In Fig. 15 we depict the correlation matrix plot of the first two eigenvalues of the nucleon two-point functions corresponding to the two states shown in Fig. 11. Here again, the second state, which may correspond to the Roper resonance, shows more correlation than the ground state.
@@
In Fig. 23 of Appendix C we show additional correlation matrix for the light nuclei. The top pane corresponds to Deuteron correlators at $m_{u}=m_{d}=m_{s}$ (left) and $m_{u}=m_{d}=m_{c}$ (right), respectively. The bottom pane represents the same for 4He. It clearly shows that correlation increases at the lighter quark mass and also increases with the atomic number. The eigenvalues corresponding to 4He correlators at the strange quark mass are correlated to a large extent, and increases further as the quark mass become lighter.
This analysis leads us to the following observations:
  1. 1.
@@
    If the statistics is high and the effective mass is reasonably good we do not find much correlation of eigenvalues across iterations. In that case fitting these eigenvalues with a constant term will be reliable and provide better estimate compared to exponential fitting. This could be a good choice in multi-operator GEVP setup, that is merging of GEVP and TGEVP together. We will explore that in future
  2. 2.
    For the higher eigenvalues, correlations are larger than that for the case of lowest eigenvalue. With higher statistics we find the correlation in the higher eigenvalues also decreases.
  3. 3.
    For multi-nucleons we find correlations to be higher than the single nucleon. This is a concern and needs to be addressed properly in a multinucleon calculation, particularly considering the presence of dense spectrum. It remains to be seen how this correlation evolves with higher statistics in these systems.
@@
Figure 14: Correlation matrix plot corresponding to the second eigenvalue of the two-point correlation function of $1^{-}$ state in charmonia ($J/\psi(2S)$ state, corresponding to the first excited state of the top write plot of Fig. 20